# Creating Automated Quantum-Assisted Solutions for Optimization Problems

### v1

@@ I Introduction
Many industries face complex optimization problems such as production planning, vehicle routing and many more. Classical algorithms typically either require prohibiting computation times or deliver suboptimal solutions. Quantum computing (QC) has the prospect to reduce the number of computational steps needed or improve the solution quality in the future. The development of tools to take on QC optimization is therefore of great interest to a wide range of industrial applications and would help integrating them into the arsenal of computation tools already available to solve them. However, venturing into quantum-assisted solutions requires a heavy investment for industries as specialists from various domains are required for the formulation, decomposition, encoding and algorithm tuning. Running quick tests with state-of-the-art methods rarely leads to satisfying results due to the complexity of the matter and the experimental stage of noisy quantum computing, which contrasts to the often miraculous claims by the hardware and software providers.
@@
The research project “Quantum-enhanced Services & Tools for Industrial Applications” (QuaST) attempts to make QC-assisted solutions more accessible to end users and allow researchers to test their methods on relevant application cases. As a central result of QuaST, we propose an integrated software tool that bundles existing knowledge in a modular and flexible way. The concept and groundwork has been laid out in [poggel_recommending_2023]. The underlying structure of the automation engine is a directed acyclic graph of options that are weighed against each other, called the QuaST decision tree (QDT). Its goal is to bring QC to the end users. In return, it also allows researchers to connect with practitioners from logistics, finance and the many other industry areas where optimization problems are relevant.
@@
  • the concept for the expansion of its features.
@@
The prototypical implementation we provide is written in Python and relies heavily on Qiskit [javadi-abhari_quantum_2024]. It guides the user through the process of solving an optimization problem with quantum-enhanced methods like VQE or QAOA, starting with the generation of the problem to the selection of a simulated or real quantum backend. The capacities of the QDT are growing, both in the number of accessible options (the “size” of the QDT) and its features. A shortlist of planned features is given in section VI. End users can start their quantum journey at the point they are familiar with their application domain. Researchers can evaluate their methods in various setups, bringing quantum technology and its users closer together.
@@
The authors of this software package strive for an open-source publication and are currently in the process of reviewing the codebase, documenting it and ensuring its quality.
@@ II Related Work
Despite the growth in research around quantum-assisted methods for optimization problems (at the time of writing, an arXiv search for the terms “quantum computing optimization” in paper abstracts yields roughly 1500 papers over the last 12 months), many researchers focus on very specific questions such as algorithmic advancements or using a specific method on a standard optimization problem. With the current state of QC and the open question of quantum advantage, progress towards abstraction layers gets a lot less attention (a similar arXiv search finds around 50 papers, with many false positives). Nonetheless, some researchers have shed some light on the topic from different perspectives.
@@ III Systematic Approach to Quantum-Assisted Solutions for Optimization Problems
To begin with, automation requires a systematic understanding. For the purpose of the QuaST decision tree, a first step has been made in [poggel_recommending_2023]. There, five layers have been identified: problem formulation, encoding, algorithm selection, classical optimizer selection and compilation/hardware selection. Hardware control, compilation and transpilation have been simplified to one layer with the plan to use existing tools and efforts for good compilation passes etc. The focus of [poggel_recommending_2023] as of this paper lies on the application side of the quantum-assisted solution.
A strictly layered approach certainly has its limits as pointed out in the original paper [poggel_recommending_2023]. From the point of view of the QuaST decision tree, the main shortcoming is the focus on variational quantum algorithms (VQA) like the Quantum Approximate Optimization Algorithm (QAOA) [farhi_quantum_2014] or the Variational Quantum Eigensolver (VQE) [peruzzo_variational_2014]. These algorithms have a specific structure in how they need to be set up (in particular regarding the involvement of a classical optimization algorithm). The QuaST decision tree in its current form also includes other algorithm types, for instance for classical comparison (brute-force sampling, Tabu sampling [palubeckis_multistart_2004]). Even state-of-the-art VQAs often have a more complex relationship between classical and quantum resources (e.g. intermediate “filter” processing steps [amaro_filtering_2022] or adaptive variants [turati_benchmarking_2023]). To allow this flexibility, the QuaST decision tree does not arrange its nodes in layers, but rather connected blocks of connected nodes. In fig. 1, these are denoted as orange boxes. Blocks share the characteristic of layers that once exited, a block cannot be reentered (unless by doing a partial restart to correct some error). However, they can be arranged more freely and, with their connections, form a directed acyclic graph. On the implementation level, these blocks have a one-to-one connection to the files of the centerpiece of the QuaST decision tree, the decisiontree/interactive subpackage. Its implementation is described in detail in section IV.
@@ IV Implementation
This chapter describes the concrete implementation of the QDT and serves as an entry point and reference guide for users and developers alike. It is mainly concerned with the framework that allows the QDT to function, with its building blocks and their relation to each other. For some examples on how to use the decision tree or contribute to it, see section V. In the remainder of this section, Section IV-A describes the algorithm behind the decision tree. Section IV-B gives more detail on its building blocks: nodes, queries, builders and hyperparameters. Then, section IV-C shows how specific nodes can be automated provided they find the correct objects to work with (e.g., optimizers).
@@ IV-A Basic Algorithm
The basic algorithm of the QDT is simple to understand: It executes code from a series of nodes that are arranged as a directed acyclic graph. The nodes modify a common dictionary called problem_data. In doing so, they perform all necessary tasks and are, in principle, unlimited in what they can do. There are QDT nodes to load the problem, encode it, request input from the user and so on. For the latter, the QDT provides various query objects described in section IV-B1. The QDT package can be simply executed via python -m decisiontree whith triggers the execution of a first node (printing a welcome message) and continues with problem loading. The behavior of the nodes can be altered with a config.json file at the project root. It is used to set levels of automation, specify locations to store data, and save access tokens for proprietary hardware and software.
@@
The pseudocode of the QDT is given in algorithm 1. It describes a mere framework with any actual transformations of the problem data being left to the nodes. However, it provides a fully flexible structure to manipulate the problem data of an optimization problem in any way, to build quantum-enhanced solutions, and to interpret their results. For a more concrete example of a node, see algorithm 2 for the loading or generation of an optimization problem.
@@ V Demonstration
In spirit, the QDT follows a user-centered design approach. The two main user types are the end user and the developer. The former are expected to use the integrated, interactive version of the QDT via python -m decisiontree after which they can load their optimization problem and are guided through the actual decision tree. The latter act as an end user when exploring and benchmarking solution paths, but contribute to the QDT by expanding it.
@@ V-A The End User Workflow
The QDT is in a prototypical stage and limited in the classes of optimization problems it can handle, but will grow with time. Optimization problems are simply defined in a human-readable JSON format containing the necessary data to define the problem. A TSP instance, for example, is defined by its adjacency matrix, containing the distances between the cities. Additional information is not required, but can be given. For example, city locations simplify the plotting of a solution, or labeling the paths between the cities may provide information on the vehicle that is used. The latter is a good example of information that is outside of the scope of the TSP, but may be very relevant for the interpretation of the result from a business perspective.
The from_dict method of the problem classes specifies the JSON format for the problem instances. In the future, it will be gathered at a central location once more problem classes are accessible through the decision tree.
@@ VI Conclusion and Outlook
  • Direct problem loading: Instead of querying the user for an input file with the optimization problem instance, the decision tree should be able to load that file directly with python -m decisiontree tsp_instance.json
@@
  • Non-interactive mode: A decision tree path should be specified by a file that allows non-interactive execution of the QuaST decision tree. More concretely, python -m decisiontree -p path_file.json should execute the path specified in path_file.json.
@@
Acknowledgement
This project is supported by the Federal Ministry for Economic Affairs and Climate Action on the basis of a decision by the German Bundestag through the project Quantum-enabling Services and Tools for Industrial Applications (QuaST). QuaST aims to facilitate the access to quantum-based solutions for optimization problems and to bridge the gap between business and technology.

### v2

@@ I Introduction
Many industries face complex optimization problems such as production planning, vehicle routing and many more. Classical algorithms typically either require prohibitively long computation times or deliver suboptimal solutions. Quantum computing (QC) has the prospect to reduce the number of computational steps needed or improve the solution quality in the future. However, the technology is not yet easily accessible to end users which hinders the adoption and further development of application-specific algorithms. Tools that lower the entry barrier and provide a level of abstraction are of great interest to a wide range of industrial applications. In particular, productive QC solutions need a close integration with classical compute power. However, venturing into quantum-assisted solutions requires a heavy investment for industries as specialists from various domains are required for the formulation, decomposition, encoding and algorithm tuning. Running quick tests with state-of-the-art methods rarely leads to satisfying results due to the complexity of the matter and the experimental stage of noisy quantum computing, which contrasts with the often miraculous claims by the hardware and software providers.
@@
The research project “Quantum-enhanced Services & Tools for Industrial Applications” (QuaST) attempts to make QC-assisted solutions more accessible to end users and allow researchers to test their methods on relevant application cases. As a central result of QuaST, we propose an integrated software tool that bundles existing knowledge in a modular and flexible way, and ultimately allows to construct an application-facing abstraction layer that can act as the top layer of any QC software stack. The concept and groundwork has been laid out in [poggel_recommending_2023]. The underlying structure of the automation engine is a directed acyclic graph of options that are weighed against each other, called the QuaST decision tree (QDT). Its goal is to bring QC to the end users. In return, it also allows researchers to connect with practitioners from logistics, finance and the many other industry areas where optimization problems are relevant.
@@
  • the concept for the expansion of its features aiming for a useful application-facing abstraction layer.
@@
The prototypical implementation we provide is written in Python and relies on Qiskit [javadi-abhari_quantum_2024], but can be extended with interfaces to any other language. It guides the user through the process of solving an optimization problem with quantum-enhanced methods like VQE or QAOA, starting with the generation of the problem to the selection of a simulated or real quantum backend. The capacities of the QDT are growing, both in the number of accessible options (the “size” of the QDT) and its features. A shortlist of planned features is given in section VI. End users can start their quantum journey at the point they are familiar with their application domain. Researchers can evaluate their methods in various setups, bringing quantum technology and its users closer together.
@@ II Related Work
Despite the growth in research around quantum-assisted methods for optimization problems (at the time of writing, an arXiv search for the terms “quantum computing optimization” in paper abstracts yields roughly 1500 papers over the last 12 months), many researchers focus on very specific questions such as algorithmic advancements or using a specific method on a standard optimization problem. With the current state of QC and the open question of quantum advantage, progress towards abstraction layers gets a lot less attention (a similar arXiv search finds around 50 papers, with many false positives). In particular, abstraction layers typically concentrate on the compilation and transpilation steps of quantum circuits and omit the problem of designing the right algorithm and quantum circuit starting from an application. Nonetheless, some researchers have shed some light on the topic from different perspectives.
@@ III Systematic Approach to Quantum-Assisted Solutions for Optimization Problems
To begin with, automation requires a systematic understanding. For the purpose of the QuaST decision tree, a first step has been made in [poggel_recommending_2023]. There, five layers have been identified: problem formulation, encoding, algorithm selection, classical optimizer selection, and compilation/hardware selection. Hardware control, compilation and transpilation have been simplified to one layer with the plan to use existing tools and efforts for good compilation passes etc. The focus of [poggel_recommending_2023] as of this paper lies on the application side of the quantum-assisted solution.
A strictly layered approach certainly has its limits as pointed out in [poggel_recommending_2023]. From the point of view of the QuaST decision tree, the main shortcoming is the focus on variational quantum algorithms (VQA) like the Quantum Approximate Optimization Algorithm (QAOA) [farhi_quantum_2014] or the Variational Quantum Eigensolver (VQE) [peruzzo_variational_2014]. These algorithms have a specific structure in how they need to be set up (in particular regarding the involvement of a classical optimization algorithm). The QuaST decision tree in its current form also includes other algorithm types, for instance for classical comparison (brute-force sampling, Tabu sampling [palubeckis_multistart_2004]). Even state-of-the-art VQAs often have a more complex relationship between classical and quantum resources (e.g. intermediate “filter” processing steps [amaro_filtering_2022] or adaptive variants [turati_benchmarking_2023]). To allow this flexibility, the QuaST decision tree does not arrange its nodes in layers, but rather connected blocks of connected nodes. In fig. 1, these are denoted as orange boxes. Blocks share the characteristic of layers that once exited, a block cannot be reentered (unless by doing a partial restart to correct some error). However, they can be arranged more freely and, with their connections, form a directed acyclic graph. On the implementation level, these blocks have a one-to-one connection to the files of the centerpiece of the QuaST decision tree, the decisiontree/interactive subpackage. Its implementation is described in detail in section IV.
@@ IV Implementation
This section describes the concrete implementation of the QDT and serves as an entry point and reference guide for users and developers alike. It is mainly concerned with the framework that allows the QDT to function, with its building blocks and their relation to each other. For some examples on how to use the decision tree or contribute to it, see section V. In the remainder of this section, Section IV-A describes the algorithm behind the decision tree. Section IV-B gives more detail on its building blocks: nodes, queries, builders and hyperparameters. Then, section IV-C shows how specific nodes can be automated provided they find the correct objects to work with (e.g., optimizers).
@@ IV-A Basic Algorithm
The basic algorithm of the QDT is simple to understand: It executes code from a series of nodes that are arranged as a directed acyclic graph. The nodes modify a common dictionary called problem_data. In doing so, they perform all necessary tasks and are, in principle, unlimited in what they can do. There are QDT nodes to load the problem, encode it, request input from the user and so on. For the latter, the QDT provides various query objects described in section IV-B1. The QDT package can be simply executed via python -m decisiontree which triggers the execution of a first node (printing a welcome message) and continues with problem loading. The behavior of the nodes can be altered with a config.json file at the project root. It is used to set levels of automation, specify locations to store data, and save access tokens for proprietary hardware and software.
@@
The pseudocode of the QDT is given in algorithm 1. It describes a mere framework with any actual transformations of the problem data being left to the nodes. However, it provides a fully flexible structure to manipulate the problem data of an optimization problem in any way, to build quantum-enhanced solutions, and to interpret their results. For a more concrete example of a node, see algorithm 2 for the loading or generation of an optimization problem. While user input can be collected, it is possible to configure the QDT such that it creates and implements recommendations for, e.g., the best algorithm to run depending on problem class or instance characteristics. In this operational mode, the complexity of the decision tree is shielded from the end user, resulting in an abstraction layer from the problem instance to the algorithm.
@@ V Demonstration
In spirit, the QDT follows a user-centered design approach. The two main user types are the end user and the developer. End users are expected to use the integrated version of the QDT via python -m decisiontree after which they can load their optimization problem and are guided through the actual decision tree. At the highest degree of automation, the QDT can be configured to implement it suggestions directly, requiring minimal user input. Developers act as an end user when exploring and benchmarking solution paths, but can contribute to the QDT by expanding it. Naturally, the behavior triggered by python -m decisiontree can be seamlessly loaded into ordinary python scripts (or Jupyter notebooks) which allows for the integration in specialized workflows including some problem generation from real-world data, or manual modification of the results.
@@ V-A The End User Workflow
As an optional third ingredient, the configuration of the QDT can be changed in order to, e.g., give more power to the end user in selecting options, or choose between different strategies for encoding the constraints. The QDT is in a prototypical stage and limited in the classes of optimization problems it can handle, with full support provided for TSP and MaxCut problems as a first step.The problem instances are simply defined in a human-readable JSON format containing the necessary data to define the problem. A TSP instance, for example, is defined by its adjacency matrix, containing the distances between the cities. Additional information is not required, but can be given. For example, city locations simplify the plotting of a solution, or labeling the paths between the cities may provide information on the vehicle that is used. The latter is a good example of information that is outside of the scope of the TSP, but may be very relevant for the interpretation of the result from a business perspective.
The from_dict method of the problem classes specifies the JSON format for the problem instances. Together with example files and tutorial notebooks for the usage of the QDT, they are a part of the software’s documentation, designed for easy usage by end users. The JSON format for problem instances in principle is designed in a way that requires minimal information about the problem instance, but additional metadata can be given to facilitate the integration with other parts of the software stack or a submission system for specific use cases that takes over, e.g., the retrieval of relevant TSP instances from real-world data.
@@ VI Conclusion and Outlook
  • Interface nodes: For further modularization, the QDT will implement interface nodes that, e.g., take over the submission to a quantum hardware provider where the compilation and transpilation can be executed. In particular, these nodes extracts the required data and packages it in a compatible way.
    In conclusion, the QDT provides a recipe and a demonstration of the upper part of a QC software stack for optimization. In particular, it takes over the extremely diverse and challenging task of systematically handling input in terms of different problem classes, formulations and algorithms with a diverse distribution of the computational load on classical and quantum systems. For end users, it is a foundational building block for a full solution pipeline with quantum-assisted algorithms.